Type your answer in the box. Estimated manufacturing overhead $450,000 Estimated direct labor hours 150,000 Actual direct labor hours 180,000 Based on this information, the amount of overhead allocated to a job that used 300 direct labor hours is $ (Enter your answer as a whole number.) Read about this Do you know the answer? I know it Think so Unsure No idea
Answer #1

Predetermine overhead rate = Estimated overhead/estimated labor hour = 450000/150000 = $3 per labor hour

Overhead allocated = Direct labor hour*Overhead rate

= 300*3

Overhead allocated = $900